Project

General

Profile

20151201 » History » Revision 28

Revision 27 (Boyd, Steven, 09 December 2015 07:36) → Revision 28/33 (Boyd, Steven, 09 December 2015 07:37)

Commissioning and Ops Scheduling meeting : 8/12/2015 

 To be held in CR10 in the ATLAS Building @ 15:30 pm  

 Phone access :    ReadyTalk number  
 US 866-740-1260 
 UK 0800-496-0576 
 Swiss 0800-70-53-51 
 Italy 800182592 
 Access Code: 2155402 

 h2. Agenda: 

 * Review of tasks and activities identified at last meeting 
 ** Vacuum / Hydrogen system work : Mark Tucker 
 ** Running with Focus Coil : MOM  
 * Upcoming activities : chaired by DC 
 ** "This week's MICE Hall work plan":http://micewww.pp.rl.ac.uk/attachments/download/5313    (Paul) 
 ** Running from Sunday 13th to Thursday 17th  
 *** Is it feasible to run with an absorber? Which absorber? (Chris, All) 
 *** If not, what else do we do? FC in solenoid mode is current fallback plan. 
 ** Personal ODH alarm usage documentation and training (Paul) 
 ** MOM/DC Handover procedure (Andy) 
 ** Christmas break walkthrough plans 

 * AOB 
 ** [[WorkPlanJanuary2016|To-do list]] for Maintenance shutdown (January 2016) 

 h2. Minutes 

 Present : Paul, Celeste, Sophie, Victoria, Henry, Paolo, Andy, Chris R, Steve W, Steve B, Paul H, Josef, John G, Mark Tucker, Francois, 
 Ken Long, Colin Whyte, Melissa, John C, Yordan, Craig Macwaters, A Gallagher, Ed Overton, P. Hanlet 

 * Review: 

 Vacuum work : most cables are done now. Should be finishing work off this week and then need to get to the stage 
 where controls work.  
 Hydrogen testing : leak testing done and localised.  
 Data taking : 4 days of good beam with 1 day of ISIS downtime. Took data for TOF0 alignment. Then data with FC 100A in flip mode 
 and -100 A in flip mode.  

 * Plans: 

 ** Tracker *Tracker He system work : John will work on this. Leak is large and probably can't run with He in the US tracker. Leak is roughly 
 4 bottles of He / day and is hard to deal with without taking the patch panel apart. Can we try to gum it up around the seal of 
 the butterfly? If we decide to run next week with absorber we need to get He in the tracker volume somehow.  

 ** A *A Gallagher : Working on vacuum system 
               Taking out the SSU dump trays and taking them back to DL. Americans don't give this a high priority at the moment. 

 ** M *M Courthold : wants to do a 50 A test on the DS whilst it's cold some time this week.* 
 Colin : On that note, DL have contacted FUG with bespoke modifications to the DS PSU and FUG have agreed to the modifications. Hope to have 
 a working PS back in February. 

 ** Absorber *Absorber running discussion : this was a baseline plan. We may not be able to run with Xe in the absorber due  
 to safety and cost concerns. We would have to put 1200 L of Xe in the buffer tank. Andy looked into the  
 case for Xe - we have to get a permit to do it and doesn't think we can do it in time. Can we work around  
 the buffer tank? Yes, but it's a significant engineering project. Paul says that the only thing we 
 can do in the time available is room temperature nitrogen. John C: Nitrogen at NTP is just as bad as air. A short  
 material comparison discussion (scat-mat.pdf) showed that only Xe realistically works. 

 A couple of questions arose: 
 Q1. How much is 1200l of Xe and what is the delivery time? 
 Q2. Can we get the safety case signed off and a permit issued before the end of the week. Andy will try. 

 Current plan : let's try to plan for Xe but expect to do measurements with the FC in solenoid mode. 

 ** Christmas *Christmas walkthrough : Victoria will organise plans for Christmas walkthrough. 

 ** AOB: *AOB: January list: Paul will hunt down system owners and ask them what they want to do in January so that we can develop a plan. 

 ALH : Spreadsheets need to be filled out and given to Pierrick. 
 Henry : need to sort out the temperature monitoring of Q4 and Q5 which have been swapped over after the user period is over. We 
 can do this after the user run finishes on the 18th. 

 John : wants to flush Q5 and Q6 in January. This has to go into the January plan.